40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 17 23:12:19 2013, Manasa, Update, Green Locking, can't get IR to resonate 
    Reply  Wed Jun 19 01:45:49 2013, Manasa, Update, Green Locking, Beat frequency sweep for 3FSR findIR.pngfindIR1.png
Message ID: 8721     Entry time: Wed Jun 19 01:45:49 2013     In reply to: 8714
Author: Manasa 
Type: Update 
Category: Green Locking 
Subject: Beat frequency sweep for 3FSR 

Measurements:

1. Calibrating offset :

I measured the shift in the beat frequency while scanning through the offset. Offset stepped by 50 resulted in 1MHz shift of the beat frequency.
 

2. Anti-whitening filter for beatbox output:

I made an anti-whitening filter for the beatbox output in the ALS_BEATX_FINE_I module by inverting the whitening filters that Jamie had installed in the beatbox earlier (elog).  I have kept the old anti-whitening filter in the module as well for the time-being because the new anti-whitening filter was not as good as the old one in stabilizing the servo (large error signals and unstable ALS).

 

3. Beat frequency scan for 3FSR:

With ALS loop enabled, I did an offset sweep corresponding to 3FSR (FSR = c/2L = 3.7MHz). The loop doesn't seem to be stable enough to reduce the arm fluctuation to get a resonance for IR. Time series of scan is shown below:

findIR.png

4. No-loop and in-loop spectrum:

I measured the spectrum of the error signal (C1:ALS-BEATX_FINE_I_IN1) with ALS loop enabled and disabled. To suppress the peaks at 3.2Hz and 16.5Hz, I turned ON the corresponding filters. I have recorded the error signal spectrum with only 16.5Hz res gain filter turned ON. Turning ON res gain 3.2Hz filter kicked ETM. 
Spectrum of error signal shown below:

findIR1.png

To resolve:

1. What is wrong with the new anti-whiteing filter?

2. Why would the res gain filters kick ETM and show no noise suppression?

ELOG V3.1.3-